0.3 g of an organic compound containing C, H and O on combustion yields 0.44 g of CO2 and 0.18 g of H20, with two O atoms per molecule. Find The number of H atoms in one mole of compound.

Grade:11

1 Answers

Vikas TU
7 years ago
First write the eqn and balance it.
CxHyOz           +        (2x + y/2) O2            --------->>>>>           xCO2           +        y/2 H2O
moles of CO2 = 0.44/(12x + 32x)
moles of H2O = 0.18/9y
moles of compound = 0.3/(12x + y + 16z)
Now U Can easily use mole mole analysis and solve the eqns for x y and z.